# Active Inference AI Systems for Scientific Discovery

### v3

@@ 1.1   The Abstraction Gap
1.2   The Reasoning Gap
Future systems must balance the complementary modes of thinking and reasoning as first-class architectural principles. Thinking—or slow, iterative discovery of new patterns—demands (i) world-model agents that can explore counterfactual spaces through mental simulation [26]; (ii) curiosity-driven mechanisms that reward pattern novelty over immediate task performance; and (iii) patience parameters that prevent premature convergence. Reasoning—the fast, deterministic traversal of pattern graphs—demands (i) efficient knowledge graph architectures with learned traversal policies; (ii) neuro-symbolic stacks that maintain both continuous representations and discrete logical structures[40]; and (iii) caching mechanisms that transform expensive thinking outcomes into rapid reasoning primitives. The interplay between these modes mirrors how scientists alternate between exploratory experimentation (thinking) and theoretical derivation (reasoning) as referenced in Inset 1.
The notion that “thinking is acting in an imaginary space"—as Konrad Lorenz observed—provides a foundational principle for understanding how world models enable scientific discovery. Just as biological organisms evolved the capacity to simulate actions internally before committing physical resources, AI systems with rich world models can explore vast hypothesis spaces through mental simulation. This capability transcends mere pattern matching: it enables counterfactual reasoning, experimental design optimization, and the anticipation of empirical surprises before they manifest in costly real-world experiments. World models can serve as the substrate for this imaginary action space, encoding not just correlations but causal structures that permit intervention and manipulation. The fidelity of these mental simulations—their alignment with physical reality—determines whether the system’s thoughts translate into valid discoveries.
Scientific progress thrives on disciplined risk: venturing beyond received wisdom while remaining falsifiable. Current alignment protocols deliberately dampen exploratory behaviour, biasing models toward safe completion of well-trodden trajectories. Controlled speculation frameworks—for example, curiosity-driven reinforcement learning [46] combined with Bayesian epistemic uncertainty quantification—could allow systems to seek novel hypotheses, flag them with calibrated uncertainty, and propose targeted experiments for arbitration. Mechanisms such as self-consistency voting [64], adversarial peer review, and tool-augmented chain-of-thought audits offer additional scaffolding to keep high-variance reasoning connected to empirical reality.
A key aspect of closing the abstraction and reasoning gaps, therefore, is in developing architectures that can construct and manipulate explicit symbolic representations as dynamic objects rather than static patterns—essentially giving models the ability to retain abstract concepts like conservation laws or causal structures in working memory and actively transform them through mental experimentation [37]. This requires moving beyond current approaches that merely associate patterns to instead build compositional graph-based [5] reasoning systems where abstract principles can be instantiated, violated, and reconstructed in imaginary spaces. Future systems can thus target manipulable conceptual building blocks that can be assembled into novel configurations that have not been explicitly seen in training—enabling the kind of counterfactual reasoning that can ultimately bridge the gap between correlation and causation that defines scientific thinking.
1.3   The Reality Gap
While the abstraction and reasoning gaps constrain what AI systems can represent and manipulate internally, the reality gap addresses a more fundamental limitation: the disconnect between computational models and the real world they aim to describe. As detailed in Inset 2, the necessity of empirical feedback for scientific discovery emerges from theoretical constraints—Gödel’s incompleteness theorems and Wolfram’s computational irreducibility guarantee that no purely computational system can discover all truths about nature. Scientific progress depends not on escaping these constraints but on navigating them through continuous dialogue with reality. In practice, this dialogue is inherently multi-modal [18] and spatio-temporal [48, 27]. AI systems must maintain a physically grounded latent state, with dynamics constrained by known operators and invariances (conservation laws; symmetries such as Galilean/rotational). This structured, multi-modal view reduces sample complexity and yields representations that extrapolate across space, time, and interventions.
Empirical feedback complements formal reasoning by supplying information inaccessible to purely deductive systems, thereby expanding—rather than mechanically escaping—the set of testable scientific propositions. The interplay between formal systems and empirical validation creates a bootstrap mechanism that circumvents incompleteness and irreducibility constraints. This suggests that AI systems for discovery must be fundamentally open—not just to new data, but to surprise from reality itself. Scientific history abounds with internally coherent theories that later failed empirical tests, underscoring the indispensability of continuous validation against data. Current AI systems excel at interpolation within their training distributions but struggle with the extrapolation that defines discovery. This is exacerbated by the fact that many scientific domains are characterized by sparse, expensive data and imperfect simulators. Unlike language modeling where data is abundant, a single protein crystallography experiment might take months and cost thousands of dollars. Simulations help but introduce their own biases, further contributing to the reality gap.
The synthesis presented in Inset 2 directly informs our architecture. Thinking explores for new pockets and tests boundaries; reasoning exploits discovered regularities. World models encode provisional maps of known pockets, subject to Popper’s falsification and Kuhn’s paradigm shifts. Human steering proves essential. Humans provide non-computational insight for recognizing genuine understanding, value judgments for directing exploration, and navigation through paradigm shifts where evaluation criteria themselves transform. Humans can shape the search process by encoding domain knowledge, identifying significant anomalies, and recognizing connections that form larger frameworks. When Faraday discovered electromagnetic induction, he did not deduce it from Maxwell’s equations (which did not yet exist)—he found it through experiment. Thus, productive collaborations can implement the complete scientific method: AI generates and tests hypotheses at scale; humans provide insight and judgment and empirical feedback provides critical steering. Our architecture must therefore implement a hybrid loop: physics priors guide ML surrogates, which direct active experiments, which update our understanding in continuous iteration.
Inset 2: Necessity of Empirical Feedback for Scientific Discovery The quest for scientific discovery via computation confronts a fundamental paradox. Gödel [23] proved that formal systems are incomplete, and cannot self-consistently prove all truths contained within the system, while Wolfram [66] demonstrates that computational irreducibility pervades nature. Additionally, Penrose [50, 51] contends that human insight transcends algorithms. Yet scientific theories and computations are found to be highly effective in many cases. Insight can be gained from Wolfram’s recent comment [67]: The very presence of computational irreducibility necessarily implies that there must be pockets of computational reducibility, where at least certain things are regular and predictable. It is within these pockets of reducibility that science fundamentally lives. In most cases, these pockets cannot be deduced a priori—they require empirical discovery. This connects to Popper’s [52] falsificationism: we cannot prove we have found true reducibility, but we can discover boundaries through experiments that challenge assumptions. Empirical feedback escapes Gödel’s constraints while delineating where nature permits shortcuts. Kuhn’s analysis [36] adds temporal dynamics: Science alternates between prevailing theories within established pockets and paradigm shifts that restructure understanding. AI systems must balance exploiting known regularities with flexibility to reconceptualize when evidence demands. In other words, even if equipped with highly effective abstractions and reasoning mechanisms, AI systems may not be able to reason their way to scientific truth through pure computation; they must actively probe reality through experiments and simulations, using empirical surprises to update their world models.
Physics priors
While generative models can create visually compelling outputs, they lack physical consistency—objects appear and disappear, gravity works intermittently, and causality is merely suggested rather than enforced. Mitchell [43] states that without biases to prefer some generalizations over others, a learning system cannot make the inductive leap necessary to classify instances beyond those it has already seen. Such inductive biases or physics priors—can be built-in to ensure generated realizations obey conservation laws, maintain object permanence, and support counterfactual reasoning about physical interactions.
Causal models
The current paradigm of domain-specific foundation models—from protein language models to molecular transformers—represents significant progress in encoding domain knowledge. However, these models fundamentally learn correlational patterns rather than causal mechanisms. ChemBERTa [14] can predict molecular properties through pattern matching but cannot simulate how modifying a functional group alters reaction pathways. AlphaFold [30] predicts protein structures through evolutionary patterns but does not model the physical folding process.
Scientific discovery requires models that transcend pattern recognition to capture causal dynamics. A causal molecular model would not just recognize that certain molecular structures correlate with properties—it would explain how electron density distributions cause reactivity, and how thermodynamic gradients drive reactions. This causal understanding enables the counterfactual reasoning essential to science: predicting outcomes of novel interventions never seen in training data. This architectural choice has profound implications: foundation models scale with data and compute, but causal models scale with understanding. With accumulaion of structural data, AI models can improve at interpolation. As we refine causal mechanisms, foundation models can improve at extrapolation—the essence of scientific discovery.
2   Additional Improvements
A certain level of consensus appears to be forming in the community that incremental scaling of present architectures may not deliver the qualitative leap that scientific discovery demands. Progress hinges on removing the design constraints through concurrent advances in algorithms and architectures (such as those described above) but also by improving efficiencies via hardware-software co-design and better evaluation benchmarks.
Computational Efficiency
@@ 2   Additional Improvements
Scaling laws show that models get predictably better with more data, parameter count and test time compute, yet every small gain might come at a great expense in time and/or energy. Such brute-force optimization contrasts sharply with biological economies in which sparse, event-based spikes[20] and structural plasticity[32] deliver continual learning at milliwatt scales. Bridging the gap will demand both algorithmic frugality—latent-variable models, active-learning curricula, reversible training—and hardware co-design. State-of-the-art foundation models require months of GPU time and $>10^{25}$ FLOPs to reach acceptable performance on long-horizon benchmarks. Memory-reversible Transformers [39, 70] and curriculum training [63] have recently reduced end-to-end training costs by 30–45 %, without loss of final accuracy. Similar level of cost reductions have been reported [16] leveraging energy and power draw scheduling.
@@
The von Neumann bottleneck—shuttling tensors between distant memory and compute—now dominates energy budgets [41]. Processing-in-memory fabrics [34], spiking neuromorphic cores that exploit event sparsity, analog photonic accelerators for low-latency matrix products, quantum samplers for combinatorial sub-routines [2] could open new algorithmic spaces. Realising their potential outside of niche applications, however, will require co-design of hardware, software and algorithms and extensive community effort.
@@
3   Architecture of Active Inference AI Systems
The fundamental gaps analyzed in the preceding sections are not independent failures but symptoms of a deeper architectural mismatch between current AI systems and the requirements of scientific discovery. These insights, combined with the inherently multi-scale and multi-modal nature of scientific phenomena—from molecular interactions to emergent spatio-temporal dynamics—dictate specific architectural requirements. A system capable of genuine discovery must therefore integrate: internal models that support mental experimentation, knowledge structures that grow through thinking-reasoning cycles, and verification mechanisms that ground speculation in empirical reality. No monolithic approach can address these diverse demands; instead, we propose a modular active inference architecture where specialized components work in concert. Figure 1 illustrates this architecture, and the key components include:
@@ 3   Architecture of Active Inference AI Systems
    Reality tethering through verification layers: The verification layer partitions scientific claims into formally provable statements and empirically testable hypotheses. Mathematical derivations, algorithmic properties, and logical arguments can be decomposed into proof obligations for interactive theorem provers (Lean [44], Coq [6]), creating a growing corpus of machine-verified knowledge that future reasoning can build upon. For claims beyond formal correctness—predictions about physical phenomena, chemical reactions, or biological behaviors—the system generates targeted computational simulations and experimental protocols. This dual approach acknowledges that scientific knowledge spans from mathematical certainty to empirical contingency. Crucially, failed verifications become learning opportunities, updating the system’s confidence bounds and identifying gaps between its world model and reality.
@@
    Human-steerable orchestration: Humans excel at recognizing meaningful patterns and making creative leaps; AI can perform exhaustive search and maintaining consistency across vast knowledge spaces; Well-understood computational science tools (e.g. optimal experimental design) can execute efficient agentic actions in a reliable manner. This symbiotic relationship ensures that the system’s powerful reasoning capabilities remain tethered to meaningful scientific questions, and existing algorithms are efficiently leveraged.
@@
    Proactive exploration engines: Rather than passively responding to queries (the primary mode in which language models are used currently), these systems work persistently in the backgrond to generate hypotheses, identify gaps in knowledge, and propose experiments. Driven by uncertainty quantification and novelty detection algorithms, these engines can maintain a priority queue of open questions ranked by their potential to achieve specified goals versus resource requirements. This layer enables the system to operate across multiple time horizons—pursuing rapid experiments vs long-term research campaigns that systematically map uncharted territories in the knowledge space.
@@
It is acknowledged that while the AI system can –in principle– be operated autonomously through well-defined interfaces between components, human interaction and decisions can be expected to play a key role. The architectural principles outlined above find partial instantiation in contemporary systems, though none fully realize the complete vision of scientific intelligence. Appendix A examines some current implementations through the lens of our three-gap framework, and discusses both substantial progress and persistent limitations that illuminate the path forward.
4   Limitations
@@
This inherent ambiguity in processing experimental feedback into actionable model refinements makes human judgment indispensable, not as a temporary scaffold but as a permanent architectural component. Thus, the challenge lies not merely in designing systems that can incorporate feedback, but in creating architectures that handle the full spectrum of empirical reality, including clear confirmations, ambiguous results, systematic biases and truly novel results. Effective human-AI collaboration must therefore go beyond simple oversight. This partnership becomes especially critical when experiments and computations challenge fundamental assumptions.
Finally, it has to be emphasized that modern AI systems are already useful in their present form, and are being utilized effectively by scientific research groups across the world. However, even with future improvements, these tools bring many systemic hazards [42]: a) false positives and false negatives: spurious correlations can be mistaken for laws, while cautious priors may hide real effects, and thus rigorous uncertainty metrics and adversarial falsification must be built in; b) epistemic overconfidence: large models could shrink error bars off-distribution, demanding ensemble disagreement; c) erosion of insight and rigor: over time, there is signficiant risk of researchers losing key scientific skills; d) Cost: simulation-driven exploration can consume resources long after marginal information saturates, so schedulers must weigh value against resources; e) concept drift: equipments and sensors evolve, and thus without continual residual checks and rapid retraining, predictions may silently bias. These issues have to be continually acknowledged, recognizing and safeguards should be embeded into the scientific process.
5   Summary and Outlook
This work has argued that the path to genuine scientific discovery through AI requires more than improvements in data, compute and scaling—it demands a reimagining of the underlying architecture. Current systems lack (i) abstractions that support mechanism-level reasoning, (ii) reasoning that operates counterfactually rather than correlationally, and (iii) reality coupling– particularly across multimodal, spatio-temporal measurements— that continuously calibrates beliefs to experiments and high-fidelity simulations. Although fully resolving these challenges is a long-horizon research program, immediate progress is possible by adopting an active-inference stack that (a) learns causal, multi-modal world models for internal simulation; (b) maintains persistent, versioned knowledge graphs with uncertainty; (c) routes claims either to formal proof engines or to empirical verification; and (d) operates under human steering where ambiguity and value trade-offs dominate. The aim is a shift from pattern completion to principle discovery.

### v4

@@ 3   Unhobbling Intelligence
Computational Inefficiency
@@
Scaling laws show that models get predictably better with more data, parameter count and test time compute, yet every small gain might come at a great expense in time and/or energy. Such brute-force optimization contrasts sharply with biological economies in which sparse, event-based spikes[30] and structural plasticity[50] deliver continual learning at milliwatt scales. Bridging the gap will demand both algorithmic frugality—latent-variable models, active-learning curricula, reversible training—and hardware co-design. State-of-the-art foundation models require months of GPU time and $10^{25}$ FLOPs to reach acceptable performance on long-horizon benchmarks. Memory-reversible Transformers [62, 107] and curriculum training [97] have recently reduced end-to-end training costs by 30–45 %, without loss of final accuracy. Similar level of cost reductions have been reported [25] leveraging energy and power draw scheduling.
@@
The von Neumann bottleneck—shuttling tensors between distant memory and compute—now dominates energy budgets [64]. Processing-in-memory fabrics [53], spiking neuromorphic cores that exploit event sparsity , analog photonic accelerators for low-latency matrix products, quantum samplers for combinatorial sub-routines [2] could open presently unreachable algorithmic spaces. Realising their potential outside of niche applications, however, will require co-design of hardware, software and algorithms and extensive community effort.
@@
4   Architecture for the Era of Experience
Empirical feedback complements formal reasoning by supplying information inaccessible to purely deductive systems, thereby expanding—rather than mechanically escaping—the set of testable scientific propositions. The interplay between formal systems and empirical validation creates a bootstrap mechanism that circumvents incompleteness and irreducibility constraints. This suggests that AI systems for discovery must be fundamentally open—not just to new data, but to surprise from reality itself. Scientific history abounds with internally coherent theories that later failed empirical tests, underscoring the indispensability of continuous validation against data. Current AI systems excel at interpolation within their training distributions but struggle with the extrapolation that defines discovery. This is exacerbated by the fact that many scientific domains are characterized by sparse, expensive data and imperfect simulators. Unlike language modeling where data is abundant, a single protein crystallography experiment might take months and cost thousands of dollars. Simulations help but introduce their own biases—the “sim-to-real gap" that plagues robotics extends to all of computational science. Our architecture must therefore implement a hybrid loop: physics priors guide ML surrogates, which direct active experiments, which update our understanding in continuous iteration.
Scientific Intelligence and the ‘Bitter Lesson’ The interplay between imagination and experimental validation creates a synergistic process at the heart of scientific discovery. Thinking provides the hypothesis generation engine, while empirical feedback provides the selection pressure that refines these mental models toward truth. These ideas have also been popularized by Yann LeCun [59] in the context of general machine intelligence. True understanding requires recognizing which causal relationships remain invariant as we move between the imaginary space of simplified models and the reality of full-scale experiments. The divergence between mental simulation and empirical outcome provides the richest learning signal. Rich Sutton’s ‘bitter lesson’ [90] observes that scalable, general methods often outperform handcrafted heuristics, yet recent domain-aware models like AlphaFold illustrate that judiciously chosen scientific priors can accelerate learning within compute-intensive regimes. AlphaFold’s use of SE(3)-equivariant networks and multiple-sequence-alignment priors enabled high accuracy with $\approx 170000$ labeled structures, but success still relied on petascale compute and vast unlabeled evolutionary data rather than a dramatic reduction from trillions of examples. We propose a ‘complete lesson’: intelligent systems arise from computation plus constraints with feedback from reality. The ‘bitter lesson’ correctly observes that general learning methods eventually outperform human-engineered heuristics within a fixed computational domain. However, it implicitly assumes that all relevant information is already encoded in the data. We contend that the most critical constraint for scientific intelligence is not a clever, human-designed inductive bias, but the information-rich feedback from the physical world, a qualitatively different signal that contains latent information about causality, invariance, and the validity of the model’s core assumptions that may be absent from any tractably finite, static dataset. The key insight is knowing when to leverage existing knowledge versus when to remain agnostic: When exploring many engineering domains (drug design, materials science), encode known physics. When questioning the nature of reality (quantum gravity, consciousness), avoid premature constraints. It is worth noting that much of science and engineering relies on the former scenario, and thus the art lies in choosing the right inductive biases for the problem at hand. This does not represent a retreat from the bitter lesson but its fulfillment: learning what to search for and how to search, guided by accumulated scientific knowledge and empirical feedback.
Causal Models
The current paradigm of domain-specific foundation models—from protein language models to molecular transformers—represents significant progress in encoding domain knowledge. However, these models fundamentally learn correlational patterns rather than causal mechanisms. ChemBERTa can predict molecular properties through pattern matching but cannot simulate how modifying a functional group alters reaction pathways. AlphaFold predicts protein structures through evolutionary patterns but does not model the physical folding process.
Scientific discovery demands models that transcend pattern recognition to capture causal dynamics. A causal molecular model would not just recognize that certain molecular structures correlate with properties—it would explain how electron density distributions cause reactivity, and how thermodynamic gradients drive reactions. This causal understanding enables the counterfactual reasoning essential to science: predicting outcomes of novel interventions never seen in training data. This architectural choice has profound implications: foundation models scale with data and compute, but causal models scale with understanding. As we accumulate more structural data, foundation models improve at interpolation. As we refine causal mechanisms, foundation models improve at extrapolation—the essence of scientific discovery.
Physics priors
While generative models like Sora create visually compelling outputs, they lack physical consistency—objects appear and disappear, gravity works intermittently, and causality is merely suggested rather than enforced. Mitchell [67] states that without biases to prefer some generalizations over others, a learning system cannot make the inductive leap necessary to classify instances beyond those it has already seen. Such inductive biases or physics priors—can be built-in to ensure generated realizations obey conservation laws, maintain object permanence, and support counterfactual reasoning about physical interactions.
Active Inference AI Systems to Navigate Complex Scientific Questions
Many scientific phenomena exhibit chaotic dynamics, multiscale interactions, and emergent properties that defy reductionist analysis. Climate systems, biological networks, and turbulent flows operate across scales from molecular to planetary. Traditional ML approaches that assume smooth, well-behaved functions fail catastrophically in these domains. We need architectures that can reason across scales, identify emergent patterns, and know when deterministic prediction becomes impossible. No single formal or informal computational system can accomplish these tasks, and hence we propose an AI stack. An exemplar architecture is shown in Figure 1. Some of the components of the architecture include:
@@ 4   Architecture for the Era of Experience
    Reality tethering through verification layers: The verification layer partitions scientific claims into formally provable statements and empirically testable hypotheses. Mathematical derivations, algorithmic properties, and logical arguments can be decomposed into proof obligations for interactive theorem provers (Lean [68], Coq [8]), creating a growing corpus of machine-verified knowledge that future reasoning can build upon. For claims beyond formal correctness—predictions about physical phenomena, chemical reactions, or biological behaviors—the system generates targeted computational simulations and experimental protocols. This dual approach acknowledges that scientific knowledge spans from mathematical certainty to empirical contingency. Crucially, failed verifications become learning opportunities, updating the system’s confidence bounds and identifying gaps between its world model and reality. When formal and empirical verification diverge for instance, when a mathematically valid prediction fails experimentally, the system must flag this inconsistency for human review, as the divergence may indicate modeling assumptions that require revision rather than simple parameter updates.
@@
    Human-steerable orchestration: Humans excel at recognizing meaningful patterns and making creative leaps; AI can perform exhaustive search and maintaining consistency across vast knowledge spaces; Well-understood computational science tools (e.g. optimal experimental design) can execute efficient agentic actions in a reliable manner. This symbiotic relationship ensures that the system’s powerful reasoning capabilities remain tethered to meaningful scientific questions, and existing algorithms are efficiently leveraged. We note that optimal experimental design methods can be inadequate when conditioned on misspecified models [89]; human oversight is therefore essential not just for value alignment but for detecting when the system’s foundational assumptions require revision.
@@
    Proactive exploration engines: Rather than passively responding to queries (the primary mode in which language models are used currently), these systems work persistently in the backgrond to generate hypotheses, identify gaps in knowledge, and propose experiments. Driven by uncertainty quantification and novelty detection algorithms, these engines can maintain a priority queue of open questions ranked by their potential to achieve specified goals versus resource requirements. This layer enables the system to operate across multiple time horizons—pursuing rapid experiments vs long-term research campaigns that systematically map uncharted territories in the knowledge space. The ranking of scientific goals reflects value judgments that must ultimately derive from human priorities; the system provides estimates of expected information gain and resource costs, but the weighting between scientific impact, feasibility, and broader societal considerations remains a human responsibility. The tension between uncertainty-driven exploration and novelty-driven exploration remains an open challenge [32]; current approaches typically require domain-specific tuning rather than universal solutions.
Component Interactions
Figure 1 illustrates the flow of information between components. The base reasoning model receives queries from human users and orchestrates calls to domain foundation models, which return predictions with uncertainty estimates. These predictions update the dynamic knowledge graph, which in turn informs the verification layer’s decisions about whether claims require formal proof, computational simulation, or physical experimentation. The performance tuner (via reinforcement learning) adjusts exploration-exploitation tradeoffs based on reward signals derived from successful predictions, experimental confirmations, and human feedback on the utility of discoveries. Concretely: (i) reward signals for the RL tuner originate from prediction accuracy on held-out data, confirmation rates in downstream experiments, and explicit human ratings of hypothesis quality; (ii) the time horizons are domain-dependent. The architectural principles outlined here are intended as design guidance rather than a complete implementation specification; realizing them in practice will require substantial engineering and domain-specific adaptation.
@@
It is acknowledged that while the AI system can, in principle, be operated autonomously through well-defined interfaces between components, human interaction and decisions can be expected to play a key role. The architectural principles outlined above find partial instantiation in contemporary systems, though none fully realize the complete vision of scientific intelligence. Appendix A examines some current implementations through the lens of our three-gap framework, and discusses both substantial progress and persistent limitations that illuminate the path forward. Appendix B gives high-level mathematical constructs for key components of the above system.
Challenges of Iterative Learning and Importance of Human Interactions
@@
This inherent ambiguity in processing experimental feedback into actionable model refinements makes human judgment indispensable, not as a temporary scaffold but as a permanent architectural component. Thus, the challenge lies not merely in designing systems that can incorporate feedback, but in creating architectures that handle the full spectrum of empirical reality, including clear confirmations, ambiguous results, systematic biases and truly novel results. Effective human-AI collaboration must therefore go beyond simple oversightempirical studies demonstrate that human-AI teams can outperform either alone, but only when humans maintain genuine understanding rather than deferring uncritically to AI suggestions [5]. This partnership becomes especially critical when experiments and computations challenge fundamental assumptions.
5   Outlook